OpenStudy (anonymous):

Use the finite sequence. Write the related series. Then evaluate the series. 26, 29, 32, 35, 38, 41, 44

OpenStudy (amistre64):

you will have to try to see a pattern to develop the rule for the summation expression

OpenStudy (amistre64):

or if you dont need to use summation notation .... just place a + between all the numbers instead of a ,

OpenStudy (anonymous):

26+3n, where n is the number in the sequence so the 1st will be 29 2nd will have n =2 and thus 32

OpenStudy (amistre64):

thats a fine rule

OpenStudy (anonymous):

thats what i got so far

OpenStudy (amistre64):

then sum that form n=0 to n=?

OpenStudy (anonymous):

till where?

OpenStudy (amistre64):

well, if counting very few terms is beyond your abilities ... you can always revert to the rule you devised .... 26 + 3n = 26; when n=0 26 + 3n = 44; when n=?

OpenStudy (anonymous):

26+3(6)=44

OpenStudy (amistre64):

then n = 0 to 6 :)

OpenStudy (anonymous):

Use the finite sequence. Write the related series. Then evaluate the series. 26, 29, 32, 35, 38, 41, 44 A. 26 + 29 + 32 + 35 + 38 + 41 + 44 = 219 B. 26 + 29 + 32 + 35 + 38 + 41 + 44 = 245 C. 26 - 29 - 32 - 35 - 38 - 41 - 44 = -193 D. 26 + 29 + 32 + 35 + 38 + 41 + 44 = 201

OpenStudy (amistre64):

\[\sum_{n=0}^{6}~26+3n\]

OpenStudy (amistre64):

i see, then the summation notation is beyond the scope

OpenStudy (anonymous):

i got B

OpenStudy (anonymous):

is it correct?

OpenStudy (amistre64):

26+3(0) 26+3(1) 26+3(2) 26+3(3) 26+3(4) 26+3(5) 26+3(6) ------- 6(26) + 3(0+1+2+...+6) 156 + 3(21)

OpenStudy (anonymous):

B was correct :).

OpenStudy (amistre64):

it was :)

OpenStudy (anonymous):

OMG i got a medal :)

OpenStudy (amistre64):

26(7) was what i typoed on